OpenStudy (daenio):

The Ford Car Company is testing their new side impact bags. It has been found that 5% of the air bags do not deploy on impact. a) What is the probability that exactly 5 of next 75 airbags tested will not deploy? b) What is the expected number of defective air bags in a new batch of 5000?

OpenStudy (daenio):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

a) P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 5) = (75 C 5)*(0.05)^(5)*(1-0.05)^(75-5) P(X = 5) = (75 C 5)*(0.05)^(5)*(0.95)^(75-5) P(X = 5) = (17259390)*(0.05)^(5)*(0.95)^70 P(X = 5) = (17259390)*(0.0000003125)*(0.0275836904) P(X = 5) = 0.14877427195401 So the probability that exactly 5 of next 75 airbags tested will not deploy is roughly 0.14877427195401

jimthompson5910 (jim_thompson5910):

b) Expected value = n*p Expected value = 5000*0.05 Expected value = 250 So the expected number of defective air bags in a new batch of 5000 is 250 airbags

OpenStudy (daenio):

Wow, you're awesome man. Thank you! :D

jimthompson5910 (jim_thompson5910):

you're welcome